Nitrogen at 25 C and 1.5 bar is contained in a vessel of 0.05 m3

. The 

gas is heated at constant volume until the pressure is 4.5 bar, and then 

cooled at constant pressure back to original temperature. Calculate the 

net heat flow from the air.


1
Expert's answer
2021-06-14T08:18:34-0400

"V=const,"

"\frac{p_1}{T_2}=\frac{p_2}{T_2},\implies"

"T_2=T_1\frac{p_2}{p_1}=3T_1=894~K,"

"p=const,"

"\frac{V_1}{T_2}=\frac{V_2}{T_1},\implies"

"V_2=V_1\frac{T_1}{T_2}=0.017~m^3,"

"\Delta U=p_2V_1-p_1V_2=0.2~J."